Use the fact that :

0= x y 1

a1 b1 1

a2 b2 1

to determine the equation of the line passing through the distinct points (a1, b1) and (a2, b2), where |·| stands for det(·), the determinant. 


1
Expert's answer
2021-05-31T15:47:50-0400
"\begin{vmatrix}
 x & y & 1 \\
 a_1 & b_1 & 1 \\
a_2 & b_2 & 1
\end{vmatrix}=0"

"x\begin{vmatrix}
 b_1& 1 \\
 b_2 & 1
\end{vmatrix}-y\begin{vmatrix}
 a_1 & 1 \\
 a_2 & 1
\end{vmatrix}+1\begin{vmatrix}
 a_1 & b_1 \\
 a_2 & b_2
\end{vmatrix}=0"

"x(b_1-b_2)-y(a_1-a_2)+a_1b_2-a_2b_1=0"

The equation of the line passing through the distinct points "(a_1, b_1)" and "(a_2, b_2)" in slope-intercept form


"y=\dfrac{b_1-b_2}{a_1-a_2}x+\dfrac{a_1b_2-a_2b_1}{a_1-a_2}, a_1\not=a_2"